OpenStudy (candra):

my question is 5.1q-1.4-4.4q=-0.3q-5.8

OpenStudy (anonymous):

q=-4.4

OpenStudy (candra):

can u help?

OpenStudy (candra):

how did u get that?

OpenStudy (anonymous):

see collect the like terms...bring all terms with q to one side of the eqn

OpenStudy (candra):

i'm sorry i dont get it

OpenStudy (anonymous):

it can be rearranged as follows 5.1q - 4.4q + 0.3q = -5.8 + 1.4 this gives q = -4.4

OpenStudy (anonymous):

Hope it is clear now....

OpenStudy (candra):

thanks

OpenStudy (anonymous):

pls click the GOOD ANSWER button to close the question

OpenStudy (anonymous):

CHOOSE "GOOD ANSWER"CHOOSE "GOOD ANSWER"CHOOSE "GOOD ANSWER"CHOOSE "GOOD ANSWER"CHOOSE "GOOD ANSWER"